OpenStudy (anonymous):

Find the solution to the equation 5x + 10 =-30 ? a. x = -8 b.x =8 c.x = -4 d. x = 4

OpenStudy (texaschic101):

5x + 10 = -30 --- subtract 10 from both sides 5x = -30 - 10 5x = -40 x = -40/5 x = -8

OpenStudy (anonymous):

thank you so much

OpenStudy (texaschic101):

sure thing :)